# RESEARCH ARTICLE

### v3

@@ 4.2.2  Timbre-Agnostic Transcription Baseline
We re-implement BasicPitch (Bittner et al.,, 2022) and Onsets&Frames (Hawthorne et al.,, 2017) in PyTorch as timbre-agnostic baselines.
@@ 4.2.3  Timbre-Separated Transcription Baseline
We reproduce the method of Tanaka et al., (2020). Due to the absence of official code and unspecified hyperparameters, we set the BiLSTM hidden size to 256, STFT window length to 1024 (matching the original frequency resolution), hop size to 256 (consistent with our setup), and embedding size to 12. We replace the original loss with our $L_{\text{InfoNCE}}$ and use our AMT branch for multi-pitch estimation. The resulting timbre encoder contains 4,897,776 trainable parameters.
@@ 4.3  Overall Performance Comparison
A similar trend appears in timbre-separated transcription in Table 3: our method (“Ours”) markedly surpasses the RNN-based baseline of Tanaka et al., (2020) (“Tanaka”). Notably, during training, both RNN baselines exhibit an earlier rise in validation loss, indicating overfitting to the training set. Given their much larger parameter counts, this suggests limited generalization under data constraints and further underscores the robustness, sample efficiency, and trainability of our lightweight architecture.
@@ 4.4  Ablation Study on Core Components
Comparing rows “Ours” and “noLog” in Table 2, we find that non-negative features indeed constrain the model’s expressive capacity. We further compare against BatchNorm (row “BN”), which is used in BasicPitch. While “BN” slightly outperforms our method on BACH10, it underperforms significantly on the other two test sets. Inspired by Esaki et al., (2024), we hypothesize that our synthetic data distribution aligns more closely with BACH10, and because BatchNorm stores learned statistics for inference, it suffers from poor generalization to more divergent domains, highlighting the advantage of our normalization strategy in cross-dataset robustness.
@@ 4.5.1  Limitations of Synthetic Data
Synthetic data underperforms MusicNet due to two domain gaps. First, our synthetic notes span the full 84-note range, whereas real instruments occupy limited registers; timbre is approximately consistent only within a narrow pitch range (Duan et al.,, 2008), leading to fragmented clusters when scattered across octaves. Second, synthesizers produce static timbres lacking the dynamic variation of real performances. These limitations underscore the need for human-composed and real-recorded datasets capturing authentic timbral complexity.
@@ 4.5.3  Exploring Alternative Encoding Architectures
Encoding dimension. Increasing the embedding dimension to 16 (row “D16”) yields no significant gain. Notably, Lin et al., (2021) achieves effective separation with only 6 dimensions, suggesting that even lighter encodings may suffice.
Multi-task coupling. Prior work (Manilow et al.,, 2020; Cwitkowitz et al.,, 2024) suggests multi-task learning can be mutually beneficial. We test two coupling strategies: 1) sharing the first residual block between branches (“Share”), and 2) rescaling the timbre embeddings using corresponding frame predictions before InstanceNorm (“Rescale”). Neither improves performance and often degrades it, indicating that such coupling may not suit lightweight networks.
@@
Attention mechanisms. In light of the contributions of Transformers to timbre modeling (Wu et al.,, 2024, 2020), we explore attention (Vaswani et al.,, 2017) for its potential to incorporate global context and enhance cluster compactness (not included in the table). Given the large $N\times T$ size, we restrict ourselves to linear-complexity variants. All tested configurations significantly hurt performance. Figure 5(d) shows that self-attention densifies intra-class connections but spuriously links distinct timbre manifolds. This occurs when attention weights connect ambiguous boundary regions where embeddings from different instruments appear similar, e.g., at overlapped notes, thereby bridging otherwise separable clusters. This is particularly detrimental to spectral clustering, which relies on graph connectivity rather than metric separability. To prevent such fusion, we also evaluate a Transformer variant with residual connections on the attention output, which merely maintains baseline performance. Attention might be more effective after note-level aggregation, but that prevents end-to-end training. For lightweight networks, especially when frame-level embeddings are not well separated, such operations appear unnecessary and potentially detrimental to separability.
@@ 5.  Reproducibility
The data and code used in this paper are accessible via https://github.com/madderscientist/timbreAMT.
@@ 6.  Conclusion
Several promising directions remain for future work. First, the current reliance on empirically tuned thresholds for note creation could be replaced by learned, content-adaptive thresholds. Second, training data could be enhanced by mixing instruments in complementary frequency bands to better reflect natural orchestration. Finally, our model still requires the number of instrument classes to be specified at inference time. The matching-filter approach described at the end of Section 3.7 remains a promising direction worth further exploration.

### v4

@@ 4.2.2  Timbre-Agnostic Transcription Baseline
We re-implement BasicPitch (Bittner et al., 2022) and Onsets&Frames (Hawthorne et al., 2017) in PyTorch as timbre-agnostic baselines.
@@ 4.2.3  Timbre-Separated Transcription Baseline
We reproduce the method of Tanaka et al. (2020). Due to the absence of official code and unspecified hyperparameters, we set the BiLSTM hidden size to 256, STFT window length to 1024 (matching the original frequency resolution), hop size to 256 (consistent with our setup), and embedding size to 12. We replace the original loss with our $L_{\text{InfoNCE}}$ and use our AMT branch for multi-pitch estimation. The resulting timbre encoder contains 4,897,776 trainable parameters.
@@ 4.3  Overall Performance Comparison
A similar trend appears in timbre-separated transcription in Table 3: our method (“Ours”) markedly surpasses the RNN-based baseline of Tanaka et al. (2020) (“Tanaka”). Notably, during training, both RNN baselines exhibit an earlier rise in validation loss, indicating overfitting to the training set. Given their much larger parameter counts, this suggests limited generalization under data constraints and further underscores the robustness, sample efficiency, and trainability of our lightweight architecture.
@@ 4.4  Ablation Study on Core Components
Comparing rows “Ours” and “noLog” in Table 2, we find that non-negative features indeed constrain the model’s expressive capacity. We further compare against BatchNorm (row “BN”), which is used in BasicPitch. While “BN” slightly outperforms our method on BACH10, it underperforms significantly on the other two test sets. Inspired by Esaki et al. (2024), we hypothesize that our synthetic data distribution aligns more closely with BACH10, and because BatchNorm stores learned statistics for inference, it suffers from poor generalization to more divergent domains, highlighting the advantage of our normalization strategy in cross-dataset robustness.
@@ 4.5.1  Limitations of Synthetic Data
Synthetic data underperforms MusicNet due to two domain gaps. First, our synthetic notes span the full 84-note range, whereas real instruments occupy limited registers; timbre is approximately consistent only within a narrow pitch range (Duan et al., 2008), leading to fragmented clusters when scattered across octaves. Second, synthesizers produce static timbres lacking the dynamic variation of real performances. These limitations underscore the need for human-composed and real-recorded datasets capturing authentic timbral complexity.
@@ 4.5.3  Exploring Alternative Encoding Architectures
Encoding dimension. Increasing the embedding dimension to 16 (row “D16”) yields no significant gain. Notably, Lin et al. (2021) achieves effective separation with only 6 dimensions, suggesting that even lighter encodings may suffice.
Multi-task coupling. Prior work (Manilow et al., 2020; Cwitkowitz et al., 2024) suggests multi-task learning can be mutually beneficial. We test two coupling strategies: 1) sharing the first residual block between branches (“Share”), and 2) rescaling the timbre embeddings using corresponding frame predictions before InstanceNorm (“Rescale”). Neither improves performance and often degrades it, indicating that such coupling may not suit lightweight networks.
@@
Attention mechanisms. In light of the contributions of Transformers to timbre modeling (Wu et al., 2024, 2020), we explore attention (Vaswani et al., 2017) for its potential to incorporate global context and enhance cluster compactness (not included in the table). Given the large $N\times T$ size, we restrict ourselves to linear-complexity variants. All tested configurations significantly hurt performance. Figure 5(d) shows that self-attention densifies intra-class connections but spuriously links distinct timbre manifolds. This occurs when attention weights connect ambiguous boundary regions where embeddings from different instruments appear similar, e.g., at overlapped notes, thereby bridging otherwise separable clusters. This is particularly detrimental to spectral clustering, which relies on graph connectivity rather than metric separability. To prevent such fusion, we also evaluate a Transformer variant with residual connections on the attention output, which merely maintains baseline performance. Attention might be more effective after note-level aggregation, but that prevents end-to-end training. For lightweight networks, especially when frame-level embeddings are not well separated, such operations appear unnecessary and potentially detrimental to separability.
@@ 5.  Reproducibility
The data and code used in this paper are accessible via https://github.com/madderscientist/timbreAMT and DOI: https://doi.org/10.5281/zenodo.19229826.
@@ 6.  Conclusion
Several promising directions remain for future work. First, the current reliance on empirically tuned thresholds for note creation could be replaced by learned, content-adaptive thresholds. Second, training data could be enhanced by mixing instruments in complementary frequency bands to better reflect natural orchestration. Finally, our model still requires the number of instrument classes to be specified at inference time and is unable to handle overlapping situations. The matching-filter approach described at the end of Section 3.7 remains a promising direction worth further exploration.